OpenStudy (anonymous):

Solve the quadratic equations: (x-4)^2=20 x^2+4x-8=0 4x^2-6x+1=0 3x^2=6x+7 Can someone show me how to do these?

OpenStudy (anonymous):

(x-4)(x-4)=20 foil the left

OpenStudy (mertsj):

Take the square root of both sides.

OpenStudy (mertsj):

x-4= +- sqrt20

OpenStudy (mertsj):

x=4+-sqrt20=4+-2sqrt5

OpenStudy (mertsj):

I just showed you the first one.

OpenStudy (anonymous):

once you foil it. then subtract the 20 from both sides

OpenStudy (mertsj):

Use quadratic formula on second one.

OpenStudy (anonymous):

Im stuck on the first one. None of my answer choices is the answer you gave me.

OpenStudy (mertsj):

Also use the formula for 3 and 4

OpenStudy (mertsj):

What are your answer choices?

OpenStudy (mertsj):

Are they in decimal form or what?

OpenStudy (anonymous):

use the quadratic formula.. a(x^2) + bx + c = 0 a,b,c are constants \[-b \pm \sqrt{b ^{2}-4ac}\div2a\]

OpenStudy (anonymous):

answer choices for number one: a. x=4+- 2 SR5 b. x=+- 6 SR 5 c. x=4 +- 5 SR 2 d. x=14 e. x= +- 2 SR 6

OpenStudy (mertsj):

What answers do you think I posted for the first one? Did you even read what I wrote?

OpenStudy (mertsj):

99 Mertsj x=4+-sqrt20=4+-2sqrt5

OpenStudy (anonymous):

You know, yahl can band me from this site if yahl want, but your seriously a feather. You have the worst attitude EVER!

OpenStudy (anonymous):

* B I T C H, not feather.